### v1

@@ 4 Structure of the Proof
Theorem 1.1 is based on three technical results that capture the core ideas behind universal Diophantine constructions: Coding, Lucas Sequences, and relation combining techniques. The dependencies between these results are illustrated in Figure 1.
The central challenge in obtaining a universal pair is reducing a general Diophantine equation — potentially involving many unknowns — to an equivalent equation with a fixed number of unknowns. This reduction is carried out in the aforementioned steps. In the process, the original equation is first transformed into intermediate expressions, which are then shown to be equivalent to a fixed Diophantine polynomial.
@@
[…]
Figure 1: Structure of the proof of Theorem 1.1. Multivariate polynomials are not displayed, they are used throughout.
@@ 5 Coding
If the $z_{i}$ are the given unknowns of the polynomial $P$ then $c$ consolidates them in a single large number. It is possible to define a number $\alpha_{P}$ such that the product $K=(1+c)^{\deg P}\cdot\alpha_{P}$ contains the value $P(z_{1},\ldots,z_{\nu})$ in a specific location of its base $\mathscr{B}$ expansion. [^1] More precisely, such $\alpha_{P}$ exists if $P$ only contains nonnegative coefficients; the statement can be generalized to arbitrary $P$ by adding an additional correcting term (cf. [1, Lemma 1.8] where $\alpha_{P}$ appears as $\operatorname{coeffs}(\mathscr{B})$). One can therefore translate the statement $P(z_{1},\ldots,z_{\nu})=0$ into a statement about the base $\mathscr{B}$ expansion of $K$. Ensuring that a certain base $\mathscr{B}$ digit of $K$ is zero can then be encoded by counting carries during binary addition. In turn, this condition can be translated into a divisibility condition on a binomial coefficient [1, Lemmas 1.3 and 1.4].
To formalize codes and work with natural numbers in base $\mathscr{B}$ we make use of the AFP entry on digit expansions [2]. We also provide a formal definition of the factorial for tuples of natural numbers and formalize the multinomial theorem — these are required for working with $\alpha_{P}$.
@@
To combine the various translation steps into a single theorem a comprehensive construction is required. One iteratively defines the auxiliary variables $\mathcal{L},r,\beta,\alpha,\gamma$ and constructs polynomials $\mathfrak{b},\mathscr{B},M,N_{0},N_{1},N,\mathscr{T},\mathscr{Y}\in{\mathbb{Z}}[a,f]$ as well as $c,\mathscr{K},\mathscr{S},\mathscr{R},\mathscr{X}\in{\mathbb{Z}}[a,f,g]$. All quantities depend on a given polynomial $P\in{\mathbb{Z}}[a,z_{1},\ldots,z_{\nu}]$ with parameter $a$ and $\nu$ unknowns. In particular, $\mathscr{K}$ corresponds to $K$ given above. We skip the precise definitions here for the sake of brevity (cf. [1, Definition 2.1]).
The following statement uses the notation $[j,k):=\{n\in{\mathbb{N}}:j\leq n<k\}$.
Theorem 5.2 (Coding Theorem [1, Theorem 1]).
For $a\geq 0$ and $f\neq 0$ such that $\mathfrak{b}:=\mathfrak{b}(a,f)$ is a power of two the following are equivalent:
@@
Our formalization of this theorem employs two locales. The first locale coding_variables fixes the polynomial $P$ along with two integers $a$ and $f$. Within this locale, the polynomials $\mathfrak{b},\mathscr{X},\mathscr{Y}$, as well as the auxiliary variables mentioned above, are introduced. The second locale, coding_theorem, incorporates the assumptions of Theorem 5.2, which are also used in several lemmas leading up to its proof. This structured approach proves advantageous in later stages of the formalization, ensuring that all coding-related concepts are accessed through a unified interface.
@@
and $\delta$_pos: "$\delta$ > 0"
Within this locale we formalize two separate statements corresponding to the two directions of the equivalence in Theorem 5.2. Here we present the statement of the reverse direction. Because $a$ and $f$ are fixed in the locale, $\mathscr{X}$ becomes a univariate polynomial in $g$ in the below code.
@@
$\land$ g < 2*(int $\gamma$)*$\mathfrak{b}$^$\alpha$
$\land$ $\mathscr{Y}$ dvd (2 * nat ($\mathscr{X}$ g) choose nat ($\mathscr{X}$ g))"
@@
A significant portion of the proof of this statement involves establishing bounds between the various variables and polynomials. The inequalities written in our mathematical draft were often not accurate or they were missing crucial assumptions. It was very convenient to then work out the details in Isabelle which tracked precisely which bounds had already been established and under which side conditions — for this type of mathematics, the proof assistant lived up to its name! We made heavy use of literal facts (writing ‹b > 0› instead of using an identifier like b_gt_0) which additionally improved the presentation of the proof.
@@ 6 Lucas Sequences and the Pell Equation
The Coding Theorem introduces an exponential relation — namely that the base $\mathfrak{b}$ be a power of two — and a binomial coefficient. Both need to be converted into an equivalent polynomial condition. To do so efficiently, we rely on the solutions of a Pell equation, also known as Lucas sequences.
@@
"$\chi$ A 0 = 2" | "$\chi$ A (Suc 0) = A" |
"$\chi$ A (Suc (Suc n)) = A * ($\chi$ A (Suc n)) - ($\chi$ A n)"
We then extend this recursive definition to all $n\in{\mathbb{Z}}$ by using the symmetries of these sequences, given by $\psi_{-n}(A)=-\psi_{n}(A)$ and $\chi_{-n}(A)=\chi_{n}(A)$. With these properties, it follows directly that Equation (8) holds for the entire integer-indexed sequence.
@@
where "$\psi$_int A n = (if n $\geq$ 0 then 1 else -1) * $\psi$ A (nat (abs n))"
@@
In the course of the proofs of these statements, totaling about 2,000 lines, we formalize many auxiliary properties of Lucas sequences. This includes both a wealth of elementary facts (cf. [1, Lemma 3.4]) and less trivial properties (cf. [1, Lemmas 3.7–3.12]). All of the latter had previously been stated or sketched in the literature, but this formalization, and our associated manuscript, contain their complete proofs for the first time.
@@ A Jungle of Variables to Build a Bridge
The specfic relations which will represent the exponential conditions from the previous section (power of two and binomial coefficient) rely on a dozen auxiliary definitions. This is not the only list of this length; also the Coding Theorem 5.2 as well as the final construction in Theorem 7.1 depend on comparable jungles of variables. We give Definition 6.3 here as an example that illustrates this complexity.
It is only partially possible to assign high-level meaning to the variable definitions below. Note that if $D$ is a square, then Equation 10f corresponds to a Pell equation with $d=A^{2}-4$. Moreover, the lemma pell_yields_lucas can be applied to relation (11b) to deduce that $J$ is the value of a Lucas sequence with parameter $U^{2}V$. Many of the remaining variables serve primarily as scaffolding, ensuring the necessary conditions to apply the appropriate number-theoretic lemmas.
@@
Theorem 6.4 (Bridge Theorem [1, Theorem 2]).
@@
The slightly simplified version of Theorem 6.4 stated here illustrates the role of this theorem for the big picture. In full detail (cf. [1, Theorem 2]), it establishes the equivalence of three statements, where the third statement is a slightly modified version of Item 2.
@@ 7 Eleven Unknowns in ${\mathbb{Z}}$
In the previous sections, we have detailed multiple translation theorems. Theorem 5.2 translates between an arbitrary polynomial $P$ having a (parametric) solution and a Diophantine predicate involving exponentiation and a binomial coefficient. Theorem 6.4 translates exponentiation and the binomial coefficient into more elementary Diophantine relations. Finally, Lemma 3.2 allows us to efficiently represent precisely these relations through a single polynomial.
@@
Before we can prove our two main Theorems 1.1 and 1.3 stated in the introduction, we first combine these three statements into an intermediate result. Let a polynomial $P$ with natural number unknowns be given. Our goal is to construct a polynomial $Q$ with a fixed number of unknowns which should have a solution whenever $P$ has a solution.
We do not give the elaborate definition of $Q$ here; it specifies how the various polynomials defined previously are substituted into each other [1, Definition 5.2]. In our construction, $Q$ has nine unknowns, eight of which are integers and the remaining one is a natural number. One can then prove the following theorem:
Theorem 7.1 (9 Unknowns in ${\mathbb{Z}}$ and ${\mathbb{N}}$ [1, Theorem 3]).
@@
The proof of this theorem proceeds by combining Theorems 5.2, 6.4 and 3.2. We structure the formalization of Theorem 7.1 into two Isabelle theories: the first one contains all definitions leading up to $Q$ and the second one implements the proof of the theorem itself. When defining $Q$ we make frequent use of the poly_extract command. We introduce custom syntax to avoid having to spell out all nine unknowns repeatedly.
We formalize two versions of the statement of Theorem 7.1. In the first version, $Q$ is simply an Isabelle term. The second version presented below uses $Q$ of type int mpoly. Note the use of is_nonnegative (defined in the obvious way) which we use because $P$ has natural number unknowns. The syntax z(0 := a) is used to substitute the parameter $a$ into the first variable of P.
@@
The statement of Theorem 7.1 is close to that of Theorem 1.3, the remaining step being that $Q$ has one remaining natural number unknown which needs to be converted into integer unknowns. One can introduce additional unknowns $n_{1},n_{2},n_{3}\in{\mathbb{Z}}$ and replace $n$ with $n_{1}^{2}+n_{2}^{2}+n_{3}^{2}+n_{3}$ to obtain a polynomial $\tilde{Q}$:
@@
We can use this to prove Theorem 1.3.
Proof 7.2 (Proof of Theorem 1.3).
By the Three Squares Theorem, any $n\geq 0$ can be represented as $n_{1}^{2}+n_{2}^{2}+n_{3}^{2}+n_{3}$. Therefore, $\tilde{Q}$ has a solution whenever $n\geq 0$ and $Q$ has a solution, i.e. when (13) holds. Using Theorem 7.1, we can thus deduce that $P$ has a solution whenever $\tilde{Q}$ has a solution. The preceding construction uses exactly 11 unknowns, completing the proof of Theorem 1.3.
Using the polynomial $\tilde{Q}$ it is also possible derive a universal pair, i.e. prove our Theorem 1.1. We do not formalize this theorem because it depends on a universal pair over ${\mathbb{N}}$.
Proof 7.3 (Proof of Theorem 1.1).
Let $(\nu,\delta)_{\mathbb{N}}$ be a universal pair. Therefore, every polynomial $P$ can be represented in at most this complexity. Theorem 1.3 establishes that $\tilde{Q}$ always has eleven integer unknowns, independent of $P$. Calculating the degree of $\tilde{Q}$ now gives the function $\eta(\nu,\delta)$ as claimed [1, Appendix A]. Using the universal pair $(32,12)_{\mathbb{N}}$, one can calculate the explicit value for the degree and obtains that
$(11,9.19593\cdot 10^{53})_{\mathbb{Z}}$
is universal.
Finally, we present the formal statement of Theorem 1.3. Our formalization implements the combination of Theorem 7.1 and the Three Squares Theorem in an additional 50 lines of Isabelle code. The formal statement assumes is_diophantine_over_N_with A P, i.e. that $A\subset{\mathbb{N}}$ is a Diophantine set and represented by the polynomial $P$. Then, by definition, $a\in A$ is equivalent to $\exists z_{1},\ldots,z_{\nu}\in{\mathbb{Z}}^{\nu}.\;P(a,z_{1},\ldots,z_{\nu})=0$.

### v2

@@ 4 Structure of the Proof
Our main result Theorem˜1.1 is based on three technical statements that capture the core ideas behind universal Diophantine constructions: Coding, Lucas sequences, and relation combining techniques. The dependencies between these results are illustrated in Figure˜1.
The central challenge in obtaining a universal pair is reducing a general Diophantine equation – potentially involving many unknowns – to an equivalent equation with a fixed number of unknowns. This reduction is carried out in the three main steps mentioned above. In the process, the original equation is first transformed into intermediate expressions, which are then shown to be equivalent to a fixed Diophantine polynomial.
@@
Figure 1: Structure of the proof of Theorem˜$1.$ Multivariate polynomials are not displayed, they are used throughout.
@@ 5 Coding
If the $z_{i}$ are the given unknowns of the polynomial $P$ then $c$ consolidates them in a single large number. It is possible to define a number $\alpha_{P}$ such that the product $K=(1+c)^{\deg P}\cdot\alpha_{P}$ contains the value $P(z_{1},\ldots,z_{\nu})$ in a specific location of its base $\mathscr{B}$ expansion. [^2] More precisely, such $\alpha_{P}$ exists if $P$ only contains nonnegative coefficients; the statement can be generalized to arbitrary $P$ by adding an additional correcting term (cf. [2, Lemma 1.8] where $\alpha_{P}$ appears as $\operatorname{coeffs}(\mathscr{B})$). One can therefore translate the statement $P(z_{1},\ldots,z_{\nu})=0$ into a statement about the base $\mathscr{B}$ expansion of $K$. Ensuring that a certain base $\mathscr{B}$ digit of $K$ is zero can then be encoded by counting carries during binary addition. In turn, this condition can be translated into a divisibility condition on a binomial coefficient [2, Lemmas 1.3 and 1.4].
To formalize codes and work with natural numbers in base $\mathscr{B}$ we make use of our own AFP entry on digit expansions [3]. We also provide a formal definition of the factorial for tuples of natural numbers and formalize the multinomial theorem – these are required for working with $\alpha_{P}$.
@@
To combine the various translation steps into a single theorem a comprehensive construction is required. One iteratively defines the auxiliary variables $\mathcal{L},r,\beta,\alpha,\gamma$ and constructs polynomials $\mathfrak{b},\mathscr{B},M,N_{0},N_{1},N,\mathscr{T},\mathscr{Y}\in{\mathbb{Z}}[a,f]$ as well as $c,\mathscr{K},\mathscr{S},\mathscr{R},\mathscr{X}\in{\mathbb{Z}}[a,f,g]$. All quantities depend on a given polynomial $P\in{\mathbb{Z}}[a,z_{1},\ldots,z_{\nu}]$ with parameter $a$ and $\nu$ unknowns. In particular, $\mathscr{K}$ corresponds to $K$ given above. We skip the precise definitions here for the sake of brevity (cf. [2, Definition 2.1]).
The following statement uses the notation $j,k):=\{n\in{\mathbb{N}}:j\leq n<k\}$.
Theorem 5.2 (Coding Theorem [2, Theorem 1]).
For $a,f\in{\mathbb{N}}$ and $f>0$ such that $\mathfrak{b}:=\mathfrak{b}(a,f)$ is a power of two, the following are equivalent:
@@
Our formalization of this theorem employs two locales. The first locale coding_variables fixes the polynomial $P$ along with two integers $a$ and $f$, and defines $\delta$ to be the total degree of $P$. Within this locale, the polynomials $\mathfrak{b},\mathscr{X},\mathscr{Y}$, as well as the auxiliary variables mentioned above, are introduced. The second locale, coding_theorem, incorporates the assumptions of Theorem˜5.2, which are also used in several lemmas leading up to its proof. This structured approach proves advantageous in later stages of the formalization, ensuring that all coding-related concepts are accessed through a unified interface.
Listing 9: Coding_Theorem/Coding_Theorem.thy (line 30).
@@
and $\delta$_pos: $\delta$ > 0"
Within this locale we formalize two separate statements corresponding to the two directions of the equivalence in Theorem˜5.2. Here we present the statement of the reverse direction. Because $a$ and $f$ are fixed in the locale, $\mathscr{X}$ now becomes a univariate polynomial in $g$.
Listing 10: Coding_Theorem/Coding_Theorem.thy (line 946).
@@
$\land$ g < 2 * $\gamma$ * $\mathfrak{b}$$\alpha$
$\land$ $\mathscr{Y}$ dvd (2 * $\mathscr{X}$ g choose $\mathscr{X}$ g)"
@@
A significant portion of the proof of this statement involves establishing bounds between the various variables and polynomials. The inequalities originally stated in our mathematical draft were often not accurate or missing crucial assumptions. It was convenient to then work out the details in Isabelle which tracked precisely which bounds had already been established and under which conditions – for this type of mathematics, the proof assistant lived up to its name! We made heavy use of literal facts (writing ‹b > 0› instead of using an identifier like b_gt_0) which additionally improved the presentation of the proof (cf. the file Lower_Bounds.thy).
@@ 6 Lucas Sequences and the Pell Equation
The Coding Theorem introduces an exponential relation – namely that the base $\mathfrak{b}$ be a power of two – and a binomial coefficient. Both need to be converted into an equivalent polynomial condition. To do so efficiently, we rely on the solutions of a Pell equation, also known as Lucas sequences.
@@
Listing 11: Lucas_Sequences/Lucas_Sequences.thy (line 7).
@@
$\chi$ A 0 = 2" | $\chi$ A (Suc 0) = A" |
$\chi$ A (Suc (Suc n)) = A * ($\chi$ A (Suc n)) - ($\chi$ A n)"
We then extend these recursive definitions to all $n\in{\mathbb{Z}}$ by using the symmetries of these sequences, given by $\psi_{-n}(A)=-\psi_{n}(A)$ and $\chi_{-n}(A)=\chi_{n}(A)$. With these properties, it follows directly that Equation (8) holds for the entire integer-indexed sequence.
Listing 12: Lucas_Sequences/DFI_square_1.thy (line 386).
@@
where $\psi$_int A n = (if n $\geq$ 0 then 1 else -1) * $\psi$ A (nat (abs n))"
@@
In the course of the proofs of these statements, totaling about 2,000 lines, we formalize many auxiliary properties of Lucas sequences. This includes both a wealth of elementary facts (cf. [2, Lemma 3.4]) and less trivial properties (cf. [2, Lemmas 3.7–3.12]). All of the latter had previously been stated or sketched in the literature, but this formalization, and our associated manuscript, contain their complete proofs for the first time.
@@ A Jungle of Variables to Build a Bridge
The specific relations which will represent the exponential conditions from the previous section (power of two and binomial coefficient) rely on a dozen auxiliary definitions. This is not the only list of this length; also the Coding Theorem 5.2 as well as the final construction in Theorem 7.1 depend on comparable jungles of variables. We give Definition 6.3 here as an example that illustrates this complexity.
It is only partially possible to assign high-level meaning to the variable definitions below. Note that if $D$ is a square, then Equation˜10f corresponds to a Pell equation with $d=A^{2}-4$. Moreover, the lemma pell_yields_lucas can be applied to relation (11b) to deduce that $J$ is the value of a Lucas sequence with parameter $U^{2}V$. Many of the remaining variables serve primarily as scaffolding, ensuring the necessary conditions to apply the appropriate number-theoretic lemmas.
@@
Theorem 6.4 (Bridge Theorem [2, Theorem 2]).
@@
The slightly simplified version of Theorem 6.4 stated here illustrates the role of this theorem for the big picture. In full detail (cf. [2, Theorem 2]), it establishes the equivalence of three statements, where the third statement is a slightly modified version of Item˜2.
@@
Listing 14: Bridge_Theorem/Bridge_Theorem.thy (line 14).
@@ 7 Eleven Unknowns in $\mathbb{Z}$}
In the previous sections, we have detailed multiple translation theorems. Theorem˜5.2 translates between an arbitrary polynomial $P$ having a (parametric) solution and a Diophantine predicate involving exponentiation and a binomial coefficient. Theorem˜6.4 translates exponentiation and the binomial coefficient into more elementary Diophantine relations. Finally, Lemma˜3.3 allows us to efficiently represent precisely these relations through a single polynomial.
@@
Before we can prove our two main Theorems 1.1 and 1.3 stated in the introduction, we first combine the previous statements into an intermediate result. Let a polynomial $P$ with natural number unknowns be given. Our goal is to construct a polynomial $Q$ with a fixed number of unknowns which should have a solution whenever $P$ has a solution.
We do not give the elaborate definition of $Q$ here; it specifies how the various polynomials defined previously are substituted into each other [2, Definition 5.2]. In our construction, $Q$ has nine unknowns, eight of which are integers and the remaining one is a natural number. One can then prove the following theorem:
Theorem 7.1 (9 Unknowns in $\mathbb{Z}$} and $\mathbb{N}$} [2, Theorem 3]).
@@
The proof of this theorem proceeds by combining Theorems˜5.2, 6.4 and 3.3. We structure the formalization of Theorem˜7.1 into two Isabelle theories: the first one contains all definitions leading up to $Q$ and the second one implements the proof of the theorem itself. When defining $Q$ we make frequent use of the poly_extract command.
We formalize two versions of the statement of Theorem˜7.1. In the first version, $Q$ is simply an Isabelle term. The second version presented below uses $Q$ of type int mpoly. Note the use of is_nonnegative (defined in the obvious way) which we use because $P$ has natural number unknowns. The syntax z(0 := a) is used to substitute the parameter $a$ into the first variable of P.
Listing 15: Nine_Unknowns_N_Z/Nine_Unknowns_N_Z.thy (line 507).
@@
The statement of Theorem˜7.1 is close to that of Theorem˜1.3, the remaining step being that $Q$ still has one natural number unknown which needs to be converted into integer unknowns. One can introduce additional unknowns $n_{1},n_{2},n_{3}\in{\mathbb{Z}}$ and replace $n$ with $n_{1}^{2}+n_{2}^{2}+n_{3}^{2}+n_{3}$ to obtain a polynomial $\tilde{Q}$:
@@
We can use this to prove Theorem˜1.3.
Proof 7.2 (Proof of Theorem˜1.3).
By the three squares Theorem, any $n\geq 0$ can be represented as $n_{1}^{2}+n_{2}^{2}+n_{3}^{2}+n_{3}$. Therefore, $\tilde{Q}$ has a solution whenever $n\geq 0$ and $Q$ has a solution, i.e. when (13) holds. Using Theorem˜7.1, we can thus deduce that $P$ has a solution whenever $\tilde{Q}$ has a solution. The preceding construction uses exactly 11 unknowns, completing the proof of Theorem˜1.3.
To formally prove Theorem˜1.3, our formalization implements the combination of Theorem˜7.1 and the three squares theorem in an additional 50 lines of Isabelle code. The formal statement assumes is_diophantine_over_N_with A P, that is, $A\subset{\mathbb{N}}$ is a Diophantine set and represented by the polynomial $P$. Then, by definition, $a\in A$ is equivalent to $\exists z_{1},\ldots,z_{\nu}\in{\mathbb{Z}}^{\nu}.\;P(a,z_{1},\ldots,z_{\nu})=0$.
Listing 16: Eleven_Unknowns_Z.thy (line 168).
@@
Using the polynomial $\tilde{Q}$ it is also possible to derive a universal pair, i.e. prove our Theorem˜1.1.
Proof 7.3 (Proof of Theorem˜1.1).
Let $(\nu,\delta)_{\mathbb{N}}$ be a universal pair. Therefore, every polynomial $P$ can be represented in at most this complexity. Theorem 1.3 establishes that $\tilde{Q}$ always has eleven integer unknowns, independent of $P$. Calculating the degree of $\tilde{Q}$ now gives the function $\eta(\nu,\delta)$ as claimed [2, Section 6]. Assuming the universal pairs $(58,4)_{\mathbb{N}}$ and $(32,12)_{\mathbb{N}}$, one can calculate the explicit value of $\eta$ and obtains that
$(11,1.68105\cdot 10^{63})\qquad\text{and}\qquad(11,9.50818\cdot 10^{53})$
are universal.
To formally prove Theorem˜1.1, we invoke our tool poly_degree Q_tilde_poly and combine the result of the automated computation with some manual evaluation and simplification. This ultimately computes the function $\eta(\nu,\delta)$ which we define in Isabelle exactly as defined in Theorem˜1.1.
The definition of universal pairs over the natural numbers is formalized as follows. Mirroring our mathematical notation, we introduce the custom syntax ($\nu$, $\delta$)N in Isabelle.
Listing 17: Universal_Pairs.thy (line 5).
definition universal_pair_N ("(_, _)N") where
[…]
is_diophantine_over_N_with A P))"
An analogous definition is made for universal_pair_Z, which introduces the custom syntax ($\nu$, $\delta$)Z.
At last, Theorem˜1.1 can be stated concisely.
Listing 18: Universal_Pairs.thy (line 15).
theorem universal_pairs_Z_from_N: "($\nu$, $\delta$)N $\Longrightarrow$ (11, $\eta$ $\nu$ $\delta$)Z"
The specific universal pairs stated above are now just special cases of this theorem. For example, the second of the two pairs (the smaller one) reads:
Listing 19: Universal_Pairs.thy (line 76).
theorem universal_pair_2:
assumes "(32, 12)N"
shows "(11, 950817549694171759711025515571236610412597656252821888)Z"